OpenStudy (anonymous):

Find the equation of the line specified. The slope is 6, and it passes through ( -4, 4). a. y = 6x + 4 c. y = 12x + 28 b. y = 6x - 20 d. y = 6x + 28

OpenStudy (mertsj):

You are given a POINT and the SLOPE. Do you think it might be a good idea to use POINT-SLOPE form?

OpenStudy (luigi0210):

And if you need the formula.. \[\LARGE y-y_{1}=m(x-x_{1})\] m=slope \[\large (x_{1},~y_{1})\]

OpenStudy (luigi0210):

Then afterwards get it into \[\LARGE y=mx+b\]

OpenStudy (anonymous):

whats the answer im on a time crunch

OpenStudy (luigi0210):

If you read the CoC it states we are here "To give help not answers" so.. sorry. :/ But this stuff is pretty simple and straight-forward.. just stop being a lazy student >_>

OpenStudy (anonymous):

meani

OpenStudy (adrynicoleb):

They are right though...It says it in the CoC and in the terms and conditions. ._. This is a study site, not a site where people give out answers. Helping, not cheating.